Susan delivers 108 newspapers in 90 minutes. Stan delivers the same number of newspaper 120 minutes. which of the following formula would allow you to determine the time required to deliver the newspapers by both Susan and Stan working together
Answer by lwsshak3(11628)

Susan delivers 108 newspapers in 90 minutes. Stan delivers the same number of newspaper 120 minutes. which of the following formula would allow you to determine the time required to deliver the newspapers by both Susan and Stan working together
**
let x=minutes required by Susan and Stan working together to deliver the newspapers
Susan's work rate=108/90
Stan's work rate=108/120
working together Susan and Stan can deliver 108 newspapers in x minutes
108x/90+108x/120=108
LCD: 90*120
120*108x+90*108x=90*120*108
120x+90x=90*120
210x=10800
x≈51 minutes